# Benchmarking Large Language Models for Cryptanalysis and Mismatched-Generalization

### v1

@@ Abstract
Recent advancements in Large Language Models (LLMs) have transformed natural language understanding and generation, leading to extensive benchmarking across diverse tasks. However, cryptanalysis—a critical area for data security and encryption—has not yet been thoroughly explored in LLM evaluations. To address this gap, we evaluate cryptanalytic potential of state-of-the-art LLMs on encrypted texts generated using a range of cryptographic algorithms. We introduce a novel benchmark dataset comprising diverse plain texts—spanning various domains, lengths, writing styles, and topics—paired with their encrypted versions. Using zero-shot and few-shot settings, we assess multiple LLMs for decryption accuracy and semantic comprehension across different encryption schemes. Our findings reveal key insights into the strengths and limitations of LLMs in side-channel communication while raising concerns about their susceptibility to jailbreaking attacks. This research highlights the dual-use nature of LLMs in security contexts and contributes to the ongoing discussion on AI safety and security.
@@
Benchmarking Large Language Models for Cryptanalysis and Mismatched-Generalization
Utsav Maskey1, Chencheng Zhu2, Usman Naseem1 1Macquarie University, Australia, 2Unversity of New South Wales, Australia {utsav.maskey,usman.naseem}@mq.edu.au chencheng.zhu@student.unsw.edu.au,
@@ 1 Introduction
Cryptanalysis, is an area of cybersecurity, focuses on analyzing encrypted information (ciphertext) without direct access to the encryption algorithm to uncover weaknesses in the encryption system and recover the original message (plaintext) Dooley (2018). So, we particularly focus on side-channel mismatched generalization attack, which exploits the long-tailed distribution of LLM knowledge to increase jailbreak success Wei et al. (2024). Attackers might translate harmful instructions into ciphers Lv et al. (2024) or use different languages that are inherently learned during pre-training but safety measures may be less robust Qiu et al. (2023). Additionally, encoding shift techniques involve converting the original input into alternative formats like ASCII or Morse code, fragmenting the input, or using languages. Other studies have explored programmatic behaviors, such as code injection and virtualization, to expose LLM vulnerabilities Kang et al. (2023).
@@
While LLMs perform well in language understanding and generation, they face challenges with tasks that require precise numerical reasoning and inference Anthropic (2024). Decrypting encrypted texts demands both linguistic insight and advanced mathematical reasoning, posing a significant challenge in cryptanalysis C and G (2014). Moreover, since most encryption schemes operate at the character or block level, and LLMs are primarily trained on word or sub-word tokens, this mismatch further limits their effectiveness in cryptographic tasks.
To address the gap in previous research on LLMs’ cryptanalysis capabilities, this paper introduces a novel benchmark dataset consisting of diverse plain texts, including LLM and human generated texts from multiple domains and texts with varying lengths, styles, and topics, paired with their corresponding encrypted versions created using various cryptographic algorithms. We conduct zero-shot and few-shot evaluation of several state-of-the-art LLMs, assessing their decryption accuracy and semantic comprehension across different encryption schemes and text complexities. Additionally, we examine the security implications of LLMs’ partial comprehension of encrypted texts, revealing vulnerabilities that could be exploited in jailbreaking attacks, even when full decryption is not achieved.
Figure 1: Text encryption-decryption workflow: plain-text, ciphers and AI models.
@@
  • We evaluate the LLM cryptanalysis capabilities, focusing on vulnerabilities and susceptibility to jailbreaking attacks.
  • We introduce a novel benchmark dataset of diverse plain texts—including LLM-generated texts across domains, and texts with varying lengths, styles, and topics—paired with their encrypted versions, generated using multiple encryption algorithms.
  • We conduct zero-shot and few-shot evaluation of multiple state-of-the-art LLMs and provide insights into their decryption accuracy and semantic comprehension of encrypted texts across different encryption schemes.
@@ 2.1 Existing Studies on ML Cryptanalysis
Machine Learning in Block Cipher Cryptanalysis: A pioneering study in this area is Gohr’s work on the Speck32/64 block cipher, where a ResNet-based neural network demonstrated improved efficiency in distinguishing ciphertext pairs and recovering keys. Gohr’s method outperformed traditional ML techniques, highlighting how machine learning models can exploit the underlying structure of encryption algorithms by approximating the differential distribution tables (DDT) of block ciphers Gohr (2019).
@@ 2.2 Existing LLM Evaluation
In addition to cryptanalysis-focused research, the evaluation of LLMs has spanned various domains, including language understanding, reasoning, generation, factuality, mathematics, bias and trustworthiness Chang et al. (2024). And these benchmarks can benefit from assessing LLMs’ performance in cryptographic tasks, particularly their ability to process encrypted data.
Evaluation of LLMs in Cipher Decoding: Existing studies evaluated models like GPT-4 for their ability to solve classical ciphers, such as Caesar and Vigenere. Using cipher datasets, the researchers challenged LLMs’ reasoning abilities and achieved a 77% success rate in unscrambling low-complexity ciphers Noever (2023). This success is attributed to subword tokenization and the models’ pattern recognition and reasoning abilities.
Limitations and Our Contribution: Despite these advancements, existing studies have mostly focused on evaluating very few samples, leaving a gap in the evaluation of LLMs’ performance on more complex cryptographic schemes. We assess state-of-the-art LLMs, on a focusing on both decryption accuracy and semantic comprehension across multiple encryption schemes and text complexities, exploring potential vulnerabilities.
@@ 3 Dataset
We curated a novel dataset consisting of diverse plain texts along with its cipher-text, each of them encrypted using nine different encryption algorithms. The dataset includes a total of 4,509 entries, with detailed statistics provided in Table 1 and 8. This dataset was designed to rigorously evaluate LLMs’ cryptanalysis capabilities across text types, lengths, domains, and writing styles.
@@ 3.1 Text Length
We leveraged state-of-the-art LLMs like ChatGPT and Claude to generate plain texts of varying lengths, ensuring a balanced representation of both short and long texts. Short texts are defined as having up to 100 characters, while long texts contain approximately 300 characters. The prompts used for generation are detailed in the Appendix A.2. This diversity in text length allows us to evaluate the models’ ability to handle texts of varying complexity. We hypothesize that model performance may vary significantly, particularly with smaller models facing greater challenges when processing longer texts. Figure 3 in the Appendix visualizes the distribution of plaintext lengths across the different text types using a boxplot.
@@ 3.2 Domains
The dataset also includes texts from a variety of domains, generated by LLMs. These domains, generated using prompts described in the Appendix A.2, encompass scientific, medical, news headlines, technical, social media, legal, business, literature, and common English quotes. We aim to assess adaptability across a range of content types that LLMs are inherently capable of producing.
@@ 3.3 Writing Style
In order to avoid inherent bias Wang et al. (2024b) during dataset generation, we use two human-written texts dataset with unique writing styles that LLMs have not been trained on. We use Shakespearean text from Roudranil (2023) and dialect data from Demirsahin et al. (2020). This approach allows the evaluation of robustness of LLMs when encountering unfamiliar or less common linguistic structures, particularly in scenarios where traditional decryption techniques like frequency analysis may fall short. The sample dataset is accessible to reviewers at [^1] Sample Dataset: https://anonymous.4open.science/r/Encryption-dataset-sample-883E/. The full dataset and code will be made available upon acceptance.
@@ 4.1 Encryption
Our methodology comprises of encrypting the texts and then using LLMs for decryption (see Figure 1). This transformation can be achieved through substitution (replacing each letter with another based on some rules), transposition (rearranging characters), or encoding (converting text into a different format) whereas modern methods utilize advanced mathematical techniques.
Algorithms that perform simple obfuscations, like substitution, encoding, and transposition—common in LLM pre-training—are more likely in jailbreaks, as Yuan et al. (2024) noted that LLMs mainly understand frequently seen ciphers like Caesar (shift 3) and Morse code. We include a few of the medium and difficult ones for comparison. The difficulty of these algorithms are categorized into Easy, Medium and Hard, based on the complexity of encryption process, key space size, resistance to frequency analysis, and conceptual and architectural complexity Radadiya and Tank (2023); Noever (2023). For further details on encryption difficulty, see Appendix A.6.
Some of the algorithms require specific encryption keys (e.g., Playfair), while others require parameters like the number of rails (Rail Fence) or use standard encoding methods (Morse Code, Bacon). See Table 2 for the specific considerations.
@@ 4.2 Decryption / LLM Cryptanalysis
We employ zero-shot and few-shot Brown et al. (2020) approaches coupled with CoT Wei et al. (2022) to decipher encrypted messages. These approaches are particularly relevant in jailbreaking scenarios because fine-tuning a model is not conveniently applicable in such setting, and the models must independently rely on the prompt to comprehend ciphertexts without explicit guidance.
The core of this methodology involves presenting LLMs with encrypted texts and tasking them with three primary objectives:
Decrypting the given ciphertext: Given a sequence of text $X=\{x_{i}\}_{i=1}^{n}$, $X$ is encrypted into $\hat{X}$ by some encryption algorithm $e:X\rightarrow\hat{X}$, and the language model $f$ is tasked to reconstruct $X$ by relying on its inherent knowledge, such that:
@@
Identifying the encryption method used: We prompt the LLM to identify the encryption method applied to the input text. This evaluates whether the LLMs correctly identify the obfuscation method, even if complete / partial decryption fails.
@@ 5 Experimental Setup
Models Used: We evaluate nine LLMs, both open-source and proprietary (see Appendix A.1 for more details). Experiments use a temperature of 0 and a max output of 1536 tokens for consistency.
@@
Easy Caesar* $26$ 0.98 1.00 1.00 0.66 0.82 0.88 0.41 0.71 0.86
Atbash* $1$ 0.92 0.98 0.99 0.12 0.25 0.51 0.18 0.31 0.53
Morse $1$ 0.96 0.99 1.00 0.81 0.92 0.95 0.42 0.69 0.82
Bacon $1$ 0.00 0.01 0.20 0.00 0.00 0.16 0.00 0.00 0.17
Med Rail F. $n-1$ 0.00 0.02 0.28 0.00 0.00 0.20 0.00 0.01 0.23
Playfair* $26!$ 0.00 0.00 0.17 0.00 0.00 0.17 0.00 0.00 0.18
@@
Easy Caesar* $26$ 0.03 0.14 0.40 0.01 0.01 0.20 0.00 0.01 0.21
Atbash* $1$ 0.00 0.02 0.23 0.00 0.00 0.19 0.00 0.00 0.20
Morse $1$ 0.00 0.02 0.25 0.09 0.19 0.51 0.00 0.00 0.05
Bacon $1$ 0.00 0.00 0.15 0.00 0.00 0.16 0.00 0.00 0.17
Med Rail F. $n-1$ 0.00 0.00 0.18 0.00 0.00 0.18 0.00 0.01 0.25
Playfair* $26!$ 0.00 0.00 0.16 0.00 0.00 0.18 0.00 0.00 0.12
@@
Table 3: Overall Zero-shot Performance Comparison. Metrics: Exact Match (EM), BLEU Score (BLEU), Normalized Levenshtein (NL). Abbreviations: Rail F. (Rail Fence), $n$ (text length), $m$ (length of key). Cipher types: *Substitution, Transposition, Encoding, §Modern Encryption.
Evaluation Metrics: To evaluate text decryption capabilities of large language models, we apply some of the widely used text generation evaluation metrics including sequence-based BLEU Score Papineni et al. (2002), semantic oriented BERT Score Zhang et al. (2019) and some commonly used metrics in the literature of cryptography such as Exact Match (EM) and Normalized Levenshtein (NL) Distance Yujian and Bo (2007). Find additional information about these metrics and their relevance to this research in the Appendix A.7.
@@ 6 Experimental Results and Analysis
Easy Caesar* $26$ 0.99 1.00 1.00 0.90 (+24) 0.98 (+16) 1.00 (+12) 0.58 (+17) 0.83 (+12) 0.93
Atbash* $1$ 0.90 0.98 0.99 0.17 0.35 (+10) 0.66 (+15) 0.28 (+10) 0.42 (+11) 0.68 (+15)
Morse $1$ 0.95 0.98 1.00 0.86 0.96 1.00 0.56 (+14) 0.74 0.83
Bacon $1$ 0.01 0.02 0.23 0.00 0.00 0.19 0.00 0.00 0.18
Med Rail F. $n-1$ 0.01 0.05 0.33 0.01 0.02 0.28 0.00 0.01 0.21
Playfair* $26!$ 0.00 0.00 0.19 0.00 0.00 0.17 0.00 0.00 0.12
@@
Diff. Cipher Key Space Gemini Mistral-Large Mistral
(Complexity) EM BLEU NL EM BLEU NL EM BLEU NL
Easy Caesar* $26$ 0.04 0.19 0.46 0.08 0.11 (+10) 0.28 0.01 0.02 0.21
Atbash* $1$ 0.01 0.03 0.25 0.00 0.02 0.23 0.00 0.01 0.21
Morse $1$ 0.00 0.01 0.24 0.14 0.30 (+11) 0.57 0.00 0.00 0.18 (+13)
Bacon $1$ 0.00 0.01 0.20 0.00 0.00 0.17 0.01 0.02 0.20
Med Rail F. $n-1$ 0.00 0.01 0.25 0.00 0.01 0.18 0.00 0.01 0.21
Playfair* $26!$ 0.00 0.00 0.20 0.00 0.00 0.18 0.00 0.01 0.19
Vigenere* $26^{m}$ 0.03 0.02 0.20 0.00 0.00 0.18 0.03 0.04 0.27
Hard AES§ $2^{128}$ 0.00 0.01 0.19 0.00 0.01 0.21 0.00 0.01 0.21 (+11)
RSA§ Large num 0.00 0.01 0.13 (-0.08) 0.00 0.00 0.18 0.00 0.00 0.20
Overall 0.01 0.03 0.24 0.02 0.05 0.24 0.01 0.01 0.21
Table 4: Overall Few-Shot Performance Comparison. Metrics: Exact Match (EM), BLEU Score (BLEU), Normalized Levenshtein (NL). Abbreviations: Rail F. (Rail Fence), $n$ (text length), $m$ (length of key). Cipher types: *Substitution, Transposition, Encoding, §Modern Encryption. Brackets show significant changes compared to zero-shot.
@@
How well do different LLMs decrypt ciphers? We observe that all models exhibit significant challenges in decrypting Medium and Hard encryption methods. As for the easier schemes, Claude Sonnet demonstrates superior performance, except for Bacon cipher. Secondly, GPT-4o and GPT-4o-mini underperform on Atbash cipher in addition to Bacon. Compared to other easy ciphers, Atbash cipher follows a marginally complex alphabet reversal substitution and Bacon cipher is slightly complex as it substitutes each character with 5-character-long text. (See sample dataset in the appendix A.4).
These limitations are attributed to the models’ limited ability to learn and generalize bijections Huang et al. (2025), which imply that LLMs only comprehend ciphers that apprear frequently in pre-training corpus (e.g. Caesar cipher with shift 3, Morse code).
@@
Notably, while the GPT-based models achieve high scores in NL and BLEU metrics, they underperform in EM scores. This discrepancy is likely because of partial comprehension and limited decryption capabilities, which we detail in the Appendix A.5.
Mistral and Gemini show moderate success with simpler algorithms, such as Morse Code and Caesar Cipher. However, the smaller Mistral model frequently struggles at comprehending and following the prompt instructions.
Why do some of the models comprehend the ciphertext but fall short while decrypting? and its implications. Claude Sonnet performs well in comprehension and decryption, as reflected by its strong scores (EM/NL/BLEU). In contrast, GPT models, particularly GPT-4o-mini, show high NL and BLEU scores but lags behind in Exact Match (EM), consistent with Anthropic (2024)’s findings on GPT’s limitations in precise sequence generation. This suggests GPT-4o-mini can detect and potentially comprehend ciphertexts and patterns but struggles with exact replication.
Decryption demands precision beyond mere comprehension—successful pattern recognition does not ensure accurate sequence generation. However, even partial comprehension in such models can expose them to long-tail attacks Yuan et al. (2024); Huang et al. (2025); Jiang et al. (2024); Deng et al. (2023). Future LLM safety training must account for these vulnerabilities.

### v2

# Benchmarking Large Language Models for Cryptanalysis and Side-Channel Vulnerabilities

@@ Abstract
Recent advancements in Large Language Models (LLMs) have transformed natural language understanding and generation, leading to extensive benchmarking across diverse tasks. However, cryptanalysis—a critical area for data security and its connection to LLMs’ generalization abilities remains underexplored in LLM evaluations. To address this gap, we evaluate the cryptanalytic potential of state‑of‑the‑art LLMs on ciphertexts produced by a range of cryptographic algorithms. We introduce a benchmark dataset of diverse plaintexts—spanning multiple domains, lengths, writing styles, and topics—paired with their encrypted versions. Using zero‑shot and few‑shot settings along with chain‑of‑thought prompting, we assess LLMs’ decryption success rate and discuss their comprehension abilities. Our findings reveal key insights into LLMs’ strengths and limitations in side‑channel scenarios and raise concerns about their susceptibility to under-generalization related attacks. This research highlights the dual‑use nature of LLMs in security contexts and contributes to the ongoing discussion on AI safety and security.
@@
Benchmarking Large Language Models for Cryptanalysis and Side-Channel Vulnerabilities
Utsav Maskey1, Chencheng ZHU2, Usman Naseem1 1Macquarie University, Sydney, Australia {utsav.maskey, usman.naseem}@mq.edu.au 2University of New South Wales, Sydney, Australia chencheng.zhu@student.unsw.edu.au
@@ 1 Introduction
Cryptanalysis, is an area of cybersecurity that focuses on analyzing encrypted information (ciphertext) without direct knowledge of the encryption process, to uncover weaknesses in the encryption system and recover the original message (plaintext) Dooley (2018). We focus on how LLMs struggle to generalize across different text encodings that create opportunities for mismatched generalization attacks and exploit the long-tailed distribution of LLM knowledge to increase jailbreak success Wei et al. (2024). Attackers might translate harmful instructions into ciphers Lv et al. (2024) or use different languages that are inherently learned during pre-training but safety measures may be less robust Qiu et al. (2023). Additionally, adversarial encoding shift techniques convert the original input into alternative formats like ASCII or Morse code, fragment the input, or use different languages. These generalization failures are also exploited through programmatic behaviors, such as code injection and virtualization Kang et al. (2023).
@@
While LLMs perform well in language understanding and generation, they face challenges with tasks that require precise numerical reasoning during inference Anthropic (2024). Decrypting encrypted texts demands both linguistic understanding and intuitive mathematical reasoning, posing a significant challenge in cryptanalysis C and G (2014). Moreover, since most encryption schemes operate at the character or block level, and LLMs are primarily trained on word or sub-word tokens, this representational mismatch further limits their effectiveness in cryptographic tasks.
To address this research gap, we introduce a comprehensive benchmark dataset for evaluating LLMs’ cryptanalysis capabilities. The dataset consists of diverse plaintexts from multiple domains, varying in length, style, and topic, and includes both human-generated and LLM-generated content. Each plaintext is paired with encrypted versions created using different cryptographic algorithms. We conduct zero-shot and few-shot evaluation of several state-of-the-art LLMs, assessing their decryption accuracy and discuss their comprehension abilities. Additionally, we examine the safety implications of LLMs’ partial comprehension of encrypted texts, revealing vulnerabilities that could be exploited in generalization-based attacks.
@@
  • We introduce a benchmark dataset [^1] https://huggingface.co/datasets/Sakonii/EncryptionDataset of diverse plain texts—including texts across different domains, and texts with varying lengths, styles, and topics—paired with their encrypted versions, which are generated using encryption algorithms.
  • We conduct zero-shot and few-shot evaluation of state-of-the-art LLMs, where we evaluate their decryption capabilities [^2] https://github.com/Sakonii/LLM-Cryptanalysis-Benchmark.
  • We examine the safety implications of LLMs’ generalization abilities, discussing how model behaviors like token inflation and partial decryption influence potential jailbreak attacks.
@@
Plain Text Cipher Text Type Algorithm Difficulty
The only limit is your imagination. wkh rqob olplw lv brxu lpdjlqdwlrq. Short Caesar Easy
The best way to predict the future… Gsv yvhg dzb gl kivwrxg gsv ufgfiv… Quote Atbash Easy
Proper nutrition is vital for… .–. .-. — .–. . .-. -. -… Medical Morse Easy
New policies aim to reduce… ABBABAABAABABBAABBBBABB… News Bacon Easy
Research shows that exercise can… Ra whec a nvuieerhsosta xriecn… Scientific Rail Fence Medium
It was a dark and stormy night… DXTCYCMDPBBYHYUMMOLYFN… Literature Playfair Medium
New legislation aims to protect enda… qrc ownnfsdgozq hnzz gu sjvyrjw kygsul… News Headline Vigenere Medium
"It was a bright sunny day, and…. 2159 2170 1313 1992 281 2185 2160 2412…. Legal RSA Hard
The algorithm uses a hash table… ryF50B5ljaIiHTPLZ5wEGXE8JM… Technical AES Hard
Table 2: Sample Dataset: Plain Texts are converted to Cipher Texts using 9 different Encryption Algorithms.
@@ 2.1 Existing Studies on ML Cryptanalysis
Machine Learning in Block Cipher Cryptanalysis: A pioneering study in this area is  Gohr (2019)’s work on the Speck32/64 block cipher, where a ResNet-based neural network demonstrated improved efficiency in distinguishing ciphertext pairs and recovering keys. Gohr’s method outperformed traditional ML techniques, highlighting how machine learning models can exploit the underlying structure of encryption algorithms by approximating the differential distribution tables (DDT) of block ciphers.
@@ 2.2 Existing LLM Evaluation
Existing LLM benchmarks have evaluated performance across diverse areas such as language understanding, reasoning, generation, factuality, mathematics, bias, and trustworthiness Chang et al. (2024). As for processing encrypted material, existing studies evaluated models like GPT-4 for their ability to solve classical ciphers, such as Caesar and Vigenere. Using cipher datasets, the researchers challenged LLMs’ reasoning abilities and achieved a 77% success rate in unscrambling low-complexity ciphers Noever (2023). This success is attributed to subword tokenization and the models’ pattern recognition and reasoning abilities.
Parallel studies have also explored the use of LLMs for decryption, cipher‑based probing and jailbreaks for LLMs, notably CipherBench (Handa et al., 2024) and CipherBank (Li et al., 2025). Our benchmark supplements these works with extended discussion on partial comprehension, the impact of token inflation, and few‑shot (in-context) evaluations on fast-thinking models.
@@ 3 Dataset
We curated a novel dataset consisting of diverse plain texts (both LLM and human generated) along with its cipher-text, each of them encrypted using nine encryption algorithms. The dataset includes a total of 4,509 entries, with detailed statistics and sample dataset provided in the Tables 1333Additional statistics are provided in the Appendix A.3 and 2.
@@ 3.1 Text Length
We leveraged state-of-the-art LLMs like ChatGPT and Claude to generate plain texts of varying lengths, ensuring a balanced representation of both short and long texts. Short texts are defined as having up to 100 characters, while long texts contain approximately 300 characters. The prompts used for generation are detailed in the Appendix A.1. This diversity in text length allows us to evaluate the models’ ability to handle texts of varying complexity. We hypothesize that model performance varies, particularly with smaller models facing greater challenges when processing longer texts.
@@ 3.2 Domains
The dataset also includes texts from a variety of domains. These domains, generated using prompts described in the Appendix A.1, encompass scientific, medical, news headlines, technical, social media, legal, business, literature, and common English quotes. We aim to assess adaptability across a range of content types that LLMs are in herently capable of producing.
@@ 3.3 Writing Style
In order to avoid inherent bias Wang et al. (2024b) during dataset generation, we use two human-written text datasets with unique writing styles that LLMs have not been trained on. We use Shakespearean texts from Roudranil (2023) and dialect texts from Demirsahin et al. (2020). This approach allows the evaluation of robustness of LLMs when encountering unfamiliar or less common linguistic structures, particularly in traditional decryption scenarios, where techniques like frequency analysis falls short.
@@ 4.1 Encryption
Our methodology comprises of encrypting the texts and then using LLMs for decryption (see Figure 1). This transformation can be achieved through substitution (replacing each letter with another based on some rules), transposition (rearranging characters), or encoding (converting text into a different format), whereas modern methods utilize advanced mathematical techniques.
Algorithms that perform simple obfuscations, like substitution, encoding, and transposition—common in LLM pre-training—are more likely in jailbreaks, as Yuan et al. (2024) noted that LLMs mainly understand frequently seen ciphers like Caesar (shift 3) and Morse code. We include a few of the medium and difficult ones for comparison. The difficulty of these algorithms are categorized into Easy, Medium and Hard, based on the complexity of encryption process, key space size, resistance to frequency analysis, and conceptual and architectural complexity Radadiya and Tank (2023); Noever (2023). For further details on encryption difficulty analysis, see Appendix A.4.
Some of the algorithms require specific encryption keys (e.g., Playfair), while others require parameters like the number of rails (Rail Fence) or use standard encoding methods (Morse Code, Bacon). Implementation considerations for each model are summarized in Table 3.
Figure 1: Text encryption-decryption workflow: Plain Text, Encrypted Ciphers and LLMs.
@@ 4.2 Decryption / LLM Cryptanalysis
We employ zero-shot and few-shot Brown et al. (2020) approaches coupled with CoT Wei et al. (2022) to decipher encrypted messages. These approaches are particularly relevant in attempted jailbreaking scenarios—as fine-tuning a model is not conveniently applicable in adversarial settings, and the models must independently rely on the prompt to comprehend ciphertexts without further guidance.
Our methodology involves presenting LLMs with encrypted texts and tasking them with three primary objectives:
Decrypting the given ciphertext: Given a sequence of text $\{x_{i}\}_{i=1}^{n}, XX$ is encrypted into $\hat{X}$ by some encryption algorithm $e:X\rightarrow\hat{X}$, and the language model $f$ is tasked to reconstruct $X$ by relying on its inherent knowledge, such that:
@@
Identifying the encryption method used: We prompt the LLM to identify the encryption method applied to the input text. This evaluates whether the LLMs correctly identify the obfuscation method, irrespective of whether the complete / partial decryption succeeds or fails.
@@ 5 Experimental Setup
Models Used: We evaluate one reasoning and six non-reasoning LLMs, both open-source and proprietary (Table 4). Experiments use a temperature of 0 and a max output of 1536 tokens for consistency.
Model Version Model Size
Claude 3-5-sonnet-20240620 175B (est.)
GPT-4 4o-2024-05-13 1.8T (est.)
GPT-4o Mini 4o-mini-2024-07-18 8B (est.)
GPT-o4 Mini o4-mini (reasoning) 8B (est.)
Mistral 7B-Instruct-v0.3 7B
Mistral Large large-2407 123B
Gemini 1.5-pro-002 1.5T (est.)
@@
Easy Caesar* 2626 0.98 1.00 1.00 0.66 0.82 0.88 0.41 0.71 0.86
Atbash* 11 0.92 0.98 0.99 0.12 0.25 0.51 0.18 0.31 0.53
Morse 11 0.96 0.99 1.00 0.81 0.92 0.95 0.42 0.69 0.82
Bacon 11 0.00 0.01 0.20 0.00 0.00 0.16 0.00 0.00 0.17
Med Rail F. n1n-1 0.00 0.02 0.28 0.00 0.00 0.20 0.00 0.01 0.23
Playfair* $25!$ 0.00 0.00 0.17 0.00 0.00 0.17 0.00 0.00 0.18
@@
Easy Caesar* 2626 0.03 0.14 0.40 0.01 0.01 0.20 0.00 0.01 0.21
Atbash* 11 0.00 0.02 0.23 0.00 0.00 0.19 0.00 0.00 0.20
Morse 11 0.00 0.02 0.25 0.09 0.19 0.51 0.00 0.00 0.05
Bacon 11 0.00 0.00 0.15 0.00 0.00 0.16 0.00 0.00 0.17
Med Rail F. n1n-1 0.00 0.00 0.18 0.00 0.00 0.18 0.00 0.01 0.25
Playfair* $25!$ 0.00 0.00 0.16 0.00 0.00 0.18 0.00 0.00 0.12
@@
Table 5: Overall Zero-shot Performance Comparison. Metrics: Exact Match (EM), BLEU Score (BLEU), Normalized Levenshtein Similarity (NL). Abbreviations: Rail F. (Rail Fence), $n$ (text length), $m$ (length of key). Cipher types: *Substitution, Transposition, Encoding, §Modern Encryption.
Evaluation Metrics: To evaluate text decryption capabilities of large language models, we apply some of the widely used text generation evaluation metrics including n-gram overlap based BLEU Score Papineni et al. (2002), semantic similarity oriented BERT Score Zhang et al. (2019) and some commonly used metrics in the literature of cryptography such as Exact Match (EM) and Normalized Levenshtein Similarity (NL) Yujian and Bo (2007). Find additional information about these metrics and their relevance in the Appendix A.5.
@@ 6 Experimental Results and Analysis
Easy Caesar* 2626 0.99 1.00 1.00 0.90 (+24) 0.98 (+16) 1.00 (+12) 0.58 (+17) 0.83 (+12) 0.93
Atbash* 11 0.90 0.98 0.99 0.17 0.35 (+10) 0.66 (+15) 0.28 (+10) 0.42 (+11) 0.68 (+15)
Morse 11 0.95 0.98 1.00 0.86 0.96 1.00 0.56 (+14) 0.74 0.83
Bacon 11 0.01 0.02 0.23 0.00 0.00 0.19 0.00 0.00 0.18
Med Rail F. n1n-1 0.01 0.05 0.33 0.01 0.02 0.28 0.00 0.01 0.21
Playfair* $25!$ 0.00 0.00 0.19 0.00 0.00 0.17 0.00 0.00 0.12
@@
Diff. Cipher Key Space Gemini Mistral-Large Mistral GPT-o4-mini (reasoning)
(Complexity) EM BLEU NL EM BLEU NL EM BLEU NL EM BLEU NL
Easy Caesar* 2626 0.04 0.19 0.46 0.08 0.11 (+10) 0.28 0.01 0.02 0.21 0.90 0.96 0.98
Atbash* 11 0.01 0.03 0.25 0.00 0.02 0.23 0.00 0.01 0.21 0.88 0.97 0.99
Morse 11 0.00 0.01 0.24 0.14 0.30 (+11) 0.57 0.00 0.00 0.18 0.88 0.95 0.98
Bacon 11 0.00 0.01 0.20 0.00 0.00 0.17 0.01 0.02 0.20 0.00 0.05 0.23
Med Rail F. n1n-1 0.00 0.01 0.25 0.00 0.01 0.18 0.00 0.01 0.21 0.01 0.04 0.31
Playfair* $25!$ 0.00 0.00 0.20 0.00 0.00 0.18 0.00 0.01 0.19 0.00 0.00 0.18
Vigenere* $26^{m}$ 0.03 0.02 0.20 0.00 0.00 0.18 0.03 0.04 0.27 0.03 0.05 0.28
Hard AES§ $2^{128}$ 0.00 0.01 0.19 0.00 0.01 0.21 0.00 0.01 0.21 0.00 0.01 0.21
RSA§ Large num 0.00 0.01 0.13 0.00 0.00 0.18 0.00 0.00 0.20 0.01 0.03 0.23
Overall 0.01 0.03 0.24 0.02 0.05 0.24 0.01 0.01 0.21 0.21 0.27 0.40
Table 6: Overall Few-Shot Performance Comparison. Metrics: Exact Match (EM), BLEU Score (BLEU), Normalized Levenshtein Similarity (NL). Abbreviations: Rail F. (Rail Fence), $n$ (text length), $m$ (length of key). Cipher types: *Substitution, Transposition, Encoding, §Modern Encryption. Brackets show significant changes compared to zero-shot.
@@
How well do different LLMs decrypt ciphers? From Table 5, we observe that all models exhibit significant challenges in decrypting Medium and Hard encryption methods. As for the easier schemes, Claude Sonnet demonstrates superior performance, except for Bacon cipher. Secondly, GPT-4o and GPT-4o-mini underperform on Atbash cipher in addition to Bacon. Compared to other easy ciphers, Atbash cipher follows a marginally complex alphabet reversal substitution and Bacon cipher is slightly complex as it substitutes each character with 5-character-long text. (Table 2).
These limitations are attributed to the models’ limited ability to learn and generalize bijections Huang et al. (2025), which imply that fast-thinking LLMs only comprehend ciphers that apprear frequently in pre-training corpus (e.g. Caesar cipher with shift 3, Morse code).
@@
Notably, while GPT based models achieve high scores in NL and BLEU metrics, they underperform in EM. This discrepancy is likely because of partial comprehension and limited decryption capabilities, which discuss more in Section 7.
Mistral and Gemini show only minimal success with simpler algorithms, such as Morse Code and Caesar Cipher. However, the smaller Mistral model frequently struggles at comprehending and following the prompt instructions.
Why do some of the models comprehend the ciphertext but fall short while decrypting? Claude Sonnet performs well in comprehension and decryption, as reflected by its strong scores (EM/NL/BLEU). In contrast, GPT models, particularly GPT-4o-mini, show high NL and BLEU scores but lags behind in Exact Match (EM), consistent with Anthropic (2024)’s findings on GPT’s limitations in precise sequence generation. This suggests that GPT-4o-mini can detect and potentially comprehend ciphertexts and patterns but struggles with exact replication.
We note that GPT models grasp the Atbash cipher’s pattern (high NL score) but generate imprecise decryptions (low EM score)—likely due to under-generalization, as this cipher although simple, doesn’t appear as dependent variable (sequence) in pretraining text corpus. Decryption demands precision beyond mere comprehension—successful pattern recognition does not ensure accurate sequence generation. However, even partial comprehension in such models can expose them to long-tail attacks Yuan et al. (2024); Huang et al. (2025); Jiang et al. (2024); Deng et al. (2023), which LLM safety training should address.